OpenStudy (anonymous):

resolve into factors y^4+2y^3z-2yz^3-z^4

OpenStudy (superhelp101):

is this algebra 1 or 2?

OpenStudy (superhelp101):

all you have to do is factor.

OpenStudy (superhelp101):

do want me to teach you how to factor or do you already know how?

OpenStudy (anonymous):

\(\Large y^4+2y^3z-2yz^3-z^4=(y^4-z^4)+(2y^3z-2yz^3)\\\Large =(y^2+z^2)(y^2-z^2)+2yz(y^2-z^2)\\\Large =(y^2-z^2)(y^2+z^2+2yz)\\\Large =(y^2-z^2)(y+z)^2\)

OpenStudy (superhelp101):

excellent work

OpenStudy (anonymous):

:) thanks